OpenStudy (anonymous):

Adam Levine wants to see if it would be worth his while to break away from Maroon 5. He hires a statistician to find his success rate with fans. The results indicated that 68 out of a 100 fans will be happy with his decision to leave Maroon 5 and go solo. Adam was hoping for a stronger show of support. Given this ratio, he wonders how many people in a 30,000 seat arena will support his decision. Help Adam determine the number of happy fans he can anticipate cheering him on.

OpenStudy (anonymous):

this cant be a test quesion is it?

OpenStudy (anonymous):

*Question

OpenStudy (anonymous):

its from an online test practice

OpenStudy (anonymous):

well i like this qustions than lol

OpenStudy (anonymous):

30,000/100=300 68*300=20,400

OpenStudy (anonymous):

hey thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

crap now i dont know how to the next problem either

OpenStudy (anonymous):

I might be able to help.

OpenStudy (anonymous):

c/2 + 4a = -2c + 3d solve for c

OpenStudy (anonymous):

I don't now either. This one is tough. Someone will probably be able to help soon though.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

He's smart.

OpenStudy (anonymous):

c/2 + 4a = -2c + 3d c/2 + 2c = -4a + 3d 2.5 c = -4a +3d c = [-4a + 3d]/2.5